OpenStudy (anonymous):

consider the region bounded by y=1/x, x=1 and x=M(M is a real number greater then 1)0 Now, calculate the volume of the solid obtained by rotating the region about the x-axis. what happens the the volume of the solid when M gets larger?

OpenStudy (anonymous):

is this a correct formula to calculate th volume?....\[\int\limits_{1}^{m}\pi(1/x)^2dx\]

OpenStudy (anonymous):

now just test values of m?

OpenStudy (anonymous):

as m gets larger the volume oviousy get larger?

OpenStudy (anonymous):

ok why shells? thanks

OpenStudy (anonymous):

v=2pi\[ 2 \Pi r h \Delta r=v\]

OpenStudy (anonymous):

i see shells give better estimate.....i must rotate around x-axis

OpenStudy (phi):

oops. I am rotating around y-axis.

OpenStudy (anonymous):

shells are just are easier that's all?

OpenStudy (phi):

so discs! pi r^2 where r= y = 1/x and volume is pi x^-2 dx

OpenStudy (phi):

which is what you did.

OpenStudy (anonymous):

cool...\[\pi \int\limits_{1}^{m}(1/x)^2 dx?....\]

OpenStudy (phi):

yes. (sorry for that detour). integrate. for what happens the the volume of the solid when M gets larger? let M-> inf and see what the limit is for the volume

OpenStudy (anonymous):

dude ...its fine im thankfull for the help iv got more if your havin fun?

OpenStudy (anonymous):

ok how do i do that??

OpenStudy (phi):

integral x^-2 ? use the power rule

OpenStudy (anonymous):

m approaches zero?

OpenStudy (anonymous):

x^-2?

OpenStudy (anonymous):

i see (1/x)^2=x^-2

OpenStudy (phi):

1/x is x^-1 and (x^-1)^2 = x^-2 integrating that is as simple as integration gets

OpenStudy (anonymous):

power rule=f prime of g(x) times g prime of x...

OpenStudy (anonymous):

so indefinite integral give limit?

OpenStudy (phi):

\[\int\limits_{1}^{m}x^{-2} dx= -x^{-1}|_{1}^{m}\]

OpenStudy (anonymous):

ok i use m!!!?]

OpenStudy (anonymous):

i plugged in m values...it increases pi/2,2pi/3....

OpenStudy (anonymous):

use the m values with out pi?

OpenStudy (phi):

you plug in the limits pi(- m^-1 - -1^-1 )= pi*(-1/m +1) or pi*(1- 1/m)

OpenStudy (anonymous):

shoot i dont get it?

OpenStudy (phi):

this is very basic calculus maybe this will help http://www.khanacademy.org/math/calculus/integral-calculus/v/introduction-to-definite-integrals

OpenStudy (phi):

\[ \pi \int\limits_{1}^{m}(1/x)^2 dx = \pi (-\frac{1}{x})|_{1}^{m} = \pi(-\frac{1}{m}+1)\]

OpenStudy (anonymous):

i did that and got....when m is 2 i get pi/2...when m is 3 i get2pi/3

OpenStudy (phi):

yes. that is right. when m is 1000 you get 999 pi/1000 you get (m-1)/m * pi as m gets big, this approaches 1*pi or pi

OpenStudy (phi):

or if you like pi*(1-1/m) = pi when m -> inf (1/very big number --> 0)

OpenStudy (phi):

do you see that the volume of this shape does not grow, but rather approaches pi as m gets bigger

OpenStudy (anonymous):

can you explain more.......i typed in 100000000000000 and got 4.4?

OpenStudy (phi):

you typed that into what?

OpenStudy (anonymous):

ti-89

OpenStudy (anonymous):

the volume really doesn't grow? its width is growing?

OpenStudy (phi):

so you are doing the integration numerically? symbolically, you get this answer http://www.wolframalpha.com/input/?i=pi+*+int+x%5E%28-2%29+dx+where+x%3D+1+to+m

OpenStudy (phi):

if you replace m with infinity in the limit, you get http://www.wolframalpha.com/input/?i=pi+*+int+x%5E%28-2%29+dx+where+x%3D+1+to+inf

OpenStudy (phi):

what is going on is the y value (1/x) is getting closer and closer to the x-axis so though the shape is getting taller,the radius y is getting smaller. the two "cancel" each other.